If (1+2x+3x 2 ) 10 = a o + a 1 x + a 2 x 2 +.......+ a 20 x 20 ,. then a 1 equals :

If (1+2x+3x2)10 = ao+ a1x + a2x2+.......+ a20x20 ,. then aequals :

Grade:11

1 Answers

Amarjeet Jayanthi
5 years ago
(1+2x+3x2)10=[(1+2x+x2)+2x2]10=[(1+x)2+(2x2)]10.
The coefficient of x in (1+2x+3x2)10 is the coefficient of x in [(1+x)2+(2x2)]10
In the expansion [(1+x)2+(2x2)]10 all the terms contain x2 except the first term.
The first term is [(1+x)2]10=(1+x)20 and the coefficient of x is 20.
Answer=20.